OpenStudy (anonymous):

Evaluate 8 to the 2 thirds power A.64 B.5 C.4 D.262,144

OpenStudy (freckles):

\[8^\frac{2}{3} =(8^\frac{1}{3})^2\] what is the cube root of 8?

OpenStudy (freckles):

or what number can you cubed that will give you 8?

OpenStudy (freckles):

\[\text{ ? }^3=8\] what goes in place of the ?

OpenStudy (anonymous):

2?

OpenStudy (freckles):

\[2^3=8 \\ \text{ so } 2=8^\frac{1}{3} \\ 8^\frac{1}{3}=2 \]

OpenStudy (freckles):

\[8^\frac{2}{3}=(8^\frac{1}{3})^2=(2)^2\] can you finish from here?

OpenStudy (anonymous):

4

OpenStudy (freckles):

right